A professor at a local university noted that the grades of her students were normally distributed with a mean of 78 and a standard deviation of 15. Overall 5 students failed the test, how many students were enrolled in her course?
Answer by stanbon(75887)

A professor at a local university noted that the grades of her students were normally distributed with a mean of 78 and a standard deviation of 15. Overall 5 students failed the test, how many students were enrolled in her course?
-----
Assuming the failure cutoff is 70%,
z(70) = (70-78)/15 = -0.5333
---
P(z < -0.333) = normalcdf(-100,-0.5333) = 0.2959
---
Let x be the number students enrolled in the course:
Solve: 0.2959x = 5 students
x = 5/0.2959
x = 16.84
Round up to 17 students in the class.
========================================
Cheers,
Stan H.
=============